OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (wasiqss):

point used does not matter, because when creating a line equation, it means all the points on the line wud b satisfied with that equation, hence any point wud gime same equation. u can check out ur self

OpenStudy (ash2326):

slope = (y2-y1)/(x2-x1) equation of line let x, y is a point on the line (y-y2)/(x-x2)=(y2-y1)/(x2-x1) x2,y2= (2,5) x1, y1=(-1,-4) (y-5)/(x-2)=(5-(-4))/(2-(-1)) (y-5)/(x-2)=9/3 (y-5)/(x-2)=3 y-5=3x-6 y=3x-1 part b No, it doesn't make any difference, whatever point one chooses, equation will come out to be the same.